OpenStudy (anonymous):

y=ax^2 + bx + c find a b and c such that the graph passes through, (-1,1) (0,1) (1,2)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

i gotta program for that ....

OpenStudy (anonymous):

16lb=_oz

OpenStudy (amistre64):

0.5x^2 + 0.5x + 1

OpenStudy (amistre64):

theres a longer complicateder way; but I just programed the HTML to do it for me..

OpenStudy (anonymous):

substitue the points in the equations , form realtionn between a,b and c. solve it u get c=1,b=.....oh! amistre gav it for ya :)

OpenStudy (amistre64):

(-1,1) (0,1) (1,2) \[(-1)^2a-1b+c=1\] \[(0)^2a+0b+c=1\] \[(1)^2a+1b+c=2\] is one way to go at it; the other is to see that you have 2 values for x that have the same y value

OpenStudy (amistre64):

so your vertex is between -1 and 0; along the x = -1/2 line

OpenStudy (amistre64):

and your y intercept is (0,1) so if anyting c=1

OpenStudy (anonymous):

using simple substitution for each equation yields: 1= a(-1)^2 +(-1)b+c = a-b+c 1=a(0)^2+(0)b + c = c 2= a(1)^2 = b(1) +c = a+b+c substitute eq. 2 in to eq.1 1=a-b+1 a=b now put eq2 and eq 1 into eq3: 2=2a+1 so that a=1/2 = b so... y=1/2x^2 +1/2x+1